OpenStudy (anonymous):

If y varies inversely as x, and y = 12 as x = 2, find y for the x-value of 6.

OpenStudy (anonymous):

A. 6 B. 5 C. 4 D. 3

OpenStudy (anonymous):

\[y =\frac{ k }{ x }\] when x=2 y=12\[12=\frac{ k }{ 2 }, k=12*2=24\]

OpenStudy (anonymous):

yeah that's not an answer either :/

OpenStudy (anonymous):

now when x=6 \[y=\frac{ 24 }{ 6 }=4\]

OpenStudy (anonymous):

oh haha that's an aswer! lol

OpenStudy (anonymous):

yep lol

OpenStudy (anonymous):

YOu're about to help me with TONS lol

OpenStudy (anonymous):

m sleepy

OpenStudy (anonymous):

sucks for you lol

OpenStudy (anonymous):

lol